OpenStudy (anonymous):

Determine which two values the following expression is between. 27.5 and 27.6 27.4 and 27.5 27.6 and 27.7 27.7 and 27.8

OpenStudy (anonymous):

\[9^{3}\sqrt{29}\]

OpenStudy (anonymous):

please help.......i'll give a medal?

OpenStudy (campbell_st):

well I don't think it lies between any of your values as you have \[9^3 \sqrt{29} = 729 \times \sqrt{29}\]

OpenStudy (ranga):

Is that 9 times cube-root of 29?

OpenStudy (anonymous):

?? what are you talking about ranga?

OpenStudy (ranga):

\[\Large 9 \times \sqrt[3]{29}\]?

OpenStudy (ranga):

\[\Large 9 \times \sqrt[3]{29} = 27.65\]